### v2

# Excitable Nonlinear Opinion Dynamics (E-NOD) for Agile Decision-Making

Abstract
We present Excitable Nonlinear Opinion Dynamics (E-NOD), which describe opinion-forming and decision-making behavior with superior “agility” in responding and adapting to fast and unpredictable changes in context, environment, or information received about available options. E-NOD derives through the introduction of a single extra term to the previously presented Nonlinear Opinion Dynamics (NOD), which have been shown to provide fast and flexible multi-agent behavior. The extra term is inspired by the fast-positive, slow-negative mixed-feedback structure of excitable systems. The agile behaviors brought about by the new excitable nature of decision-making driven by E-NOD are analyzed in a general setting and illustrated in an application to robot navigation around human movers.
I Introduction
The nonlinear opinion dynamics (NOD) presented in [1, 2] model the time evolution of opinions of a group of agents engaged in a collective decision-making process over a set of options. The derivation of NOD was tailored to model and study the principles of adaptive decision-making in biological [3, 4, 5, 6] collectives and to use these principles to design adaptive decision-making in built collectives [1, 7, 8]. From a systems perspective, NOD exhibits a mixed-feedback [9, 10] structure: opinion formation arises from the balance of a negative feedback loop that regulates the agents’ opinions to a neutral state and many positive feedback loops (both at the single-agent and network levels) that destabilize the neutral state and trigger nonlinear opinion formation.
Decision-making driven by NOD is fast and flexible [2]. It is fast because it can exhibit fast divergence from indecision even in the absence of informative inputs about the options. It is flexible because in the presence of informative inputs the sensitivity of opinion formation to the information the inputs carry is tunable and adaptive to context. Both speed and flexibility of decision-making driven by NOD are determined by the existence of a tunable threshold for opinion formation. At this threshold, negative and positive feedback are perfectly balanced and the dynamics become singular.
Figure 1: (a) Trajectory of a robot controlled with NOD (E-NOD) is shown with a blue (pink) line as it navigates towards a goal (star) in the presence of an oncoming human mover (black). The NOD robot experiences a collision; the E-NOD robot does not. (b) Opinion $z$ of the robot over time $t$. Circles denote matching time points along trajectories and opinions.
Figures 1, 4, and 5 animated in video here.
Here, we introduce a single extra term in the NOD equations that enables the following “agile” decision-making behaviors: a) autonomous sequential decision-making, not requiring any ad-hoc reset of the model state once a decision is made; b) fast “changes of mind” and adaptive responses, when the context or information about the options change rapidly and unexpectedly; c) on-demand, or event-based, opinion formation in the sense that large opinions are formed sparsely in time in the form of “decision events” and only when context requires it; d) low energy and low resource use for greedy decision-making, resulting from the dynamics being event-based. These new agile behaviors are illustrated in Fig. 1 in the context of a robot navigating around a human mover as studied in [11].
The term we introduce in NOD is a slow regulation term inspired by the dynamics of excitable (spiking) signal processing systems, widespread in nature [12, 13, 14] and foundational to spiky control systems [15] and neuromorphic engineering [16]. The result is event-based decision-making that is a form of multi-dimensional excitability in the opinion formation space, i.e., spiking can occur in any of the multiple directions corresponding to the multiple options. This multi-dimensional excitability generalizes the one-dimensional excitability of classical spiking systems. We call NOD with the extra regulation term Excitable Nonlinear Opinion Dynamics (E-NOD).
Our contributions are fivefold. First, in Section II, we present a new analysis of the singularity in NOD for a single decision-making agent and two options. We prove how a feedback gain, $K_{u}$, tunes opinion formation. We show that when $K_{u}$ gets too large, an opinion can become so robust that it will not change quickly enough if a new input arrives in favor of the alternative option (blue in Fig. 1). Second, in Section III, we present E-NOD for a single agent and two options. We show existence of the limit cycles associated with excitable behavior. We show geometrically the agility in behavior (pink in Fig. 1). Fourth, we generalize E-NOD to a network of multiple agents (Section IV). Fifth, we apply E-NOD to the social robot navigation problem (Section V).
II Fast and flexible decision-making
We recall in Section II-A the NOD [1, 2] for a single agent evolving continuously over time its real-valued opinion about two mutually exclusive options in the possible presence of input. In Section II-B we analyze stability of the neutral solution and prove conditions on feedback gain $K_{u}$ that determine the type of singularity (type of pitchfork bifurcation) in the dynamics. We show that, by shaping bifurcation branches, $K_{u}$ tunes opinion formation and enables agile decision-making. In Section II-C, we show limits on tunability of NOD, which motivates introducing E-NOD in Section III.
II-A NOD for Single Decision-Maker
We let an agent represent a single decision-maker. Let $z(t)\in\mathbb{R}$ define the agent’s opinion at time $t$ about two mutually exclusive options. The more positive (negative) is $z$, the more the agent favors (disfavors) option 1 and disfavors (favors) option 2. When $z\!=\!0$ the agent is neutral about the two options, i.e., in a state of indecision. Let $u(t)\!\geq\!0$ define the attention of the agent at time $t$ to its observations; $u$ is implemented as a gain in the dynamics. Let $b(t)\in\mathbb{R}$ define an input signal at time $t$ that represents external stimulus and/or internal bias. When $b(t)\!>\!0$ ($b(t)\!<\!0$), it provides information (evidence) in favor of option 1 (option 2).
Decision-making variables $z$ and $u$ evolve continuously over time $t$ according to the following NOD, adapted from [1, 2]:
$\tau_{z}\dot{z}$ $=-d\,z+\tanh\big{(}u\left(\alpha z\right)+b\big{)}$ (1a)
$u$ $=u_{0}+K_{u}z^{2}$ (1b)
where ${\dot{z}}:=\!dz/dt$. $\tau_{z}\!>\!0$ is a time constant, and damping coefficient $d\!>\!0$ weights the negative feedback on $z$ that regulates the neutral state $z\!=\!0$. The second term in (1a) provides a nonlinear positive feedback on $z$ with weight given by the product of $u$ and amplification coefficient $\alpha\!>\!0$, plus the effects of $b$. The saturation nonlinearity given by the $\tanh$ function enables fast-and-flexible decision-making through opinion-forming bifurcations [1, 2]. The positive feedback gain is state-dependent according to (1b) and grows with $z^{2}$. Hence, small deviations from the neutral state in response to small inputs leave attention low and do not trigger large, nonlinear opinion formation. Large enough deviations from the neutral state in response to large enough inputs cause a sharp increase in attention and trigger large, nonlinear opinion formation. The resulting implicit threshold distinguishing small and large inputs is tuned by parameters $u_{0}$ and $K_{u}$.
II-B Analysis of NOD
We study the dynamics and stability of solutions of system (1a)-(1b) using bifurcation theory. A local bifurcation refers to a change in number and/or stability of the equilibrium solutions to a nonlinear dynamical system as a (bifurcation) parameter is changed. The state and parameter values at which this change occurs is called a bifurcation point. At a bifurcation point, one or more of the eigenvalues of the Jacobian of the model must have zero real part [17, 18], i.e., a bifurcation point is a singularity of the model vector field.
We are mainly interested in the pitchfork bifurcations. There are two generic types of pitchforks. A supercritical pitchfork bifurcation describes how one stable solution becomes unstable and two stable solutions emerge as the bifurcation parameter increases. A subcritical pitchfork bifurcation describes how two unstable solutions disappear and one stable solution becomes unstable as the bifurcation parameter increases.
Our objective is to understand how thresholds of fast-and-flexible decision-making are controlled by the model parameters with the goal of designing feedback control laws for those parameters that can make decision thresholds adaptive to context. Substituting (1b) into (1a) yields
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}+K_{u}z^{2})\cdot\left(\alpha z\right)+b\Big{)}.$ (2)
We first study (2) in the case $b\!=\!0$, i.e., when there is no evidence to distinguish the options, and bifurcations are symmetric. Then, we introduce $b\!\neq\!0$ and use unfolding theory [17] to understand the effects of inputs.
Lemma 1.
(NOD Taylor expansion and singularity): Consider (2) and let $b\!=\!0$. Then the solution $z\!=\!0$ is always an equilibrium, and the Taylor expansion of (2) about $z\!=\!0$ is
$\dot{z}=\frac{1}{\tau_{z}}\Bigg{(}\left(\alpha\,u_{0}-d\right)z+\alpha\left(K_{u}-\frac{\alpha^{2}u_{0}^{3}}{3}\right)z^{3}\\ +\alpha^{3}u_{0}^{2}\left(\frac{2\alpha^{2}u_{0}^{3}}{15}-K_{u}\right)z^{5}\Bigg{)}+\mathcal{O}(z^{7}).$ (3)
A singularity exists at $(u_{0},z)\!=\!(u_{0}^{*},0)$, with $u_{0}^{*}\!=\!\frac{d}{\alpha}$. The solution $z\!=\!0$ is stable (unstable) when $u_{0}\!<\!u_{0}^{*}$ ($u_{0}\!>\!u_{0}^{*}$).
Proof.
When $z\!=\!0$ and $b\!=\!0$ the righthand side of (2) is zero so $z\!=\!0$ is always an equilibrium. We expand (2) with $b\!=\!0$ about $z\!=\!0$. The Taylor expansion of the hyperbolic tangent is $\tanh(w)\!=\!w-w^{3}/3+2w^{5}/15+\mathcal{O}(w^{7})$. Substituting this into (2) yields (3). The Jacobian $J\!=\!\frac{d\dot{z}}{dz}$ of (3) evaluated at $z\!=\!0$ is $J(0)\!=\!(\alpha\,u_{0}-d)/\tau_{z}$, which is singular when $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. When $u\!<\!u_{0}^{*}$, $J(0)\!<\!0$ so $z\!=\!0$ is exponentially stable and when $u\!>\!u_{0}^{*}$, $J(0)\!>\!0$ so $z\!=\!0$ is unstable. ∎
We next explore in Proposition 1 and Fig. 2 the effect of parameter $K_{u}$ on the cubic and quintic terms of (3) and its role in determining the type of singularity at $(u,z)\!=\!(u_{0}^{*},0)$.
Proposition 1.
($K_{u}$ determines the singularity type): Consider (2). Let $b\!=\!0$ and $u_{0}^{*}\!=\!\frac{d}{\alpha}$. The singularity of dynamics (2) at $(u,z)\!=\!(u_{0}^{*},0)$ as proved in Lemma 1 corresponds to a supercritical pitchfork bifurcation for $K_{u}\!<\!\frac{d^{3}}{3\alpha}$, a quintic pitchfork bifurcation for $K_{u}\!=\!\frac{d^{3}}{3\alpha}$, and a subcritical pitchfork bifurcation for $K_{u}\!>\!\frac{d^{3}}{3\alpha}$.
Proof.
We define cubic coefficient $p$ to be the coefficient of $\alpha z^{3}/\tau_{z}$ in (3) at $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. Then $p\!=\!(K_{u}\!-\!\frac{d^{3}}{3\alpha})$ and $p\!<\!0$ ($p\!>\!0$) for all $K_{u}\!<\!\frac{d^{3}}{3\alpha}$ ($K_{u}\!>\!\frac{d^{3}}{3\alpha}$). When $p\!<\!0$ ($p\!>\!0$), the term stabilizes (destabilizes) the system and produces a supercritical (subcritical) pitchfork bifurcation. We define quintic coefficient $q$ to be the coefficient of $(\alpha^{3}u_{0}^{2})z^{5}/\tau_{z}$ in (3) at $u_{0}\!=\!u_{0}^{*}\!=\!\frac{d}{\alpha}$. Then $q\!=\!(\frac{2d^{3}}{15\alpha}\!-\!K_{u})$. Coefficients $p$ and $q$ as functions of $K_{u}$ are plotted in Fig. 2b. The results follow by invoking the recognition problem for the quintic pitchfork [17, Prop. VI.2.14] and the form of its $Z_{2}$-symmetric universal unfolding [17, Prop. VI.3.4 and Fig. VI.3.3]. ∎
Figure 2: The effect of $K_{u}$ on the bifurcation diagram of (2) and the cubic and quintic terms of (3). (a): Bifurcation diagrams of NOD (2) with $K_{u}$ values corresponding to the vertical dotted lines in (b). Stable (unstable) solutions are shown with solid (dotted) lines. The bifurcation point is $(u_{0}^{*},0)$. (b): Coefficient $p$ ($q$) as a function of $K_{u}$ shown as a solid (dashed) black line.
Proposition 1 uncovers the key role of $K_{u}$ in tuning opinion formation: $K_{u}$ controls the supercritical versus subcritical nature of opinion formation and increasing $K_{u}$ increases opinion strength at non-neutral bifurcation branches (Fig. 2). The bifurcation diagrams in Fig. 2a are plots of the equilibrium solutions of (2), i.e., the solutions of $\dot{z}\!=\!0$, as a function of bifurcation parameter $u_{0}$, for five different values of $K_{u}$ distinguished by color. The singularity at $(u_{0},z)\!=\!(u_{0}^{*},0)$ is a pitchfork bifurcation: blue, gray, and green lines show supercritical, quintic, and subcritical solutions, respectively. The corresponding values of $K_{u}$ are indicated in Fig. 2b. As $K_{u}$ increases from 00, the supercritical pitchfork grows wider (increasing opinion strength) as $p$ grows less negative, becomes quintic when $p\!=\!0$, and becomes a subcritical pitchfork when $p\!>\!0$. For the subcritical pitchforks in Fig. 2a, two stable non-neutral solutions appear for $u_{0}\!<\!u_{0}^{*}$, through so-called saddle-node bifurcations. As $p$ grows more positive, these solutions emerge for smaller values of $u_{0}$ and increase in magnitude, reflecting increasing opinion strength.
II-C Limitation on Tuning of NOD
We prove that the region of multi-stability in the subcritical bifurcation of NOD grows as $K_{u}$ gets large.
Proposition 2.
($K_{u}$ determines region of multi-stability): Let $b=0$ and $(u_{0}^{\dagger},z^{\dagger})$ be either of the two saddle-node bifurcation points of the subcritical pitchfork of NOD (2) for $K_{u}\!>\!\frac{d^{3}}{3\alpha}$. Then $u_{0}^{\dagger}$ is a monotonically decreasing function of $K_{u}$, i.e., $\frac{\partial u_{0}^{\dagger}}{\partial K_{u}}\!<\!0$.
Proof.
Let $K_{u}^{\dagger}\!>\!\frac{d^{3}}{3\alpha}$ and $f(z,K_{u},u_{0})\!:=\!-d\,z\!+\!\tanh(\alpha z(u_{0}\!+\!K_{u}z^{2}))$. By hypothesis, $f(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})\!=\!0$. We have $\frac{\partial f}{\partial u_{0}}(z^{\dagger},K_{u}^{\dagger},u_{0}^{\dagger})=\alpha z^{\dagger}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}\!+\!K_{u}^{\dagger}(z^{\dagger})^{2}))\!\neq\!0$, since $z^{\dagger}\!\neq\!0$. Following [19], we use the implicit function theorem to show the existence of $g\!:\!\mathbb{R}^{2}\!\to\!\mathbb{R}$ such that for some neighborhood of $(z^{\dagger},K_{u}^{\dagger})$, $f(z,K_{u},g(z,K_{u}))\!=\!0$. We get:
$\frac{\partial g}{\partial K_{u}}$ $=-\left(\frac{\partial f}{\partial u_{0}}\right)^{-1}\left(\frac{\partial f}{\partial K_{u}}\right)$
$=-\frac{\alpha(z^{\dagger})^{3}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}{\alpha z^{\dagger}\tanh^{\prime}(\alpha z^{\dagger}(u_{0}^{\dagger}+K_{u}^{\dagger}(z^{\dagger})^{2}))}=-(z^{\dagger})^{2}.$
Since $z^{\dagger}\!\neq\!0$, $u_{0}^{\dagger}$ is monotonically decreasing in $K_{u}$. ∎
Proposition 2 implies that one limitation of NOD is that large $K_{u}$ can make multi-stability so large and robust that solutions can get “stuck” in one of the decision attractors unless very large inputs in favor of another decision state are applied. This is illustrated in Fig. 3A, where the first (dark blue trace) and second (light blue trace) NOD differ only in their $K_{u}$ parameters $K_{u_{1}}\!>\!K_{u_{2}}\!>\!\frac{d^{3}}{3\alpha}$ but their solutions are distinctively different. At the stimulus onset ($b\!>\!0$ for $1000\!\leq\!t\!<\!$ 100), the solution of the first NOD converges to $z\!>\!0$ much more rapidly than that of the second NOD. When the input switches values ($b\!<\!0$ for $t\!\geq\!100$), the solution of the first NOD gets stuck to a positive value, whereas the solution of the second NOD is able to track the change in input sign. This example reveals a fundamental trade-off between speed/robustness (first, dark blue NOD) and flexibility (second, light blue NOD).
Instead of aiming to fine-tune the gain $K_{u}$ around some hard to define fast/robust enough yet flexible enough decision-making behavior, we use mixed-feedback principles to make the system excitable, therefore inheriting both the speed of large-$K_{u}$ NODs and the flexibility of small-$K_{u}$ NODs. The behavior of the resulting E-NOD is shown in the third line (pink trace) of Fig. 3A. By generating “decision spikes” the E-NOD is as fast as the high $K_{u}$ NOD and as flexible as the low $K_{u}$ NOD. The rest of the paper is devoted to presenting the E-NOD model, its analysis, and its multi-agent generalization.
Figure 3: Opinion solutions of NOD and E-NOD over time and associated bifurcation diagrams. (A): Trajectories of NOD (2) and E-NOD (4) for initial condition $(z,u_{0})|_{t=0}=(0.01,0.9)$ for two $K_{u}$ values. Input signal $b$ is also shown over time. (B): Bifurcation diagrams of (2) for the two values of $K_{u}$, showing the solutions in A. moving from initial state to steady state.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) $u_{0}=0.1,b=0$
(b) $u_{0}=0.5,b=0$
(c) $u_{0}=0.7,b=0$
(d) $u_{0}=0.7,b=0.1$
Figure 4: The system solutions and its $(u_{s},z)$ phase portrait as the basal attention $u_{0}$ increases. For all plots, $d\!=\!1,\alpha\!=\!2,$ (thus $u_{0}^{*}=0.5$)$,K_{u}\!=\!2,K_{u_{s}}\!=\!6,\tau_{z}\!=\!2,\tau_{u_{s}}\!=\!20$. (Top): Example solutions of $u_{s}$ and $z$ over time, with initial condition as $(u_{s},z)|_{t=0}=(0.01,0.01)$ and additive Gaussian distributed white noise. (Bottom): The $u_{s}$-nullcline ($z$-nullcline) is shown as a blue (pink) line. Solid (dotted) lines indicate stable (unstable) branches of the $z$-nullcline with respect to (4a). Gray arrows denote the vector field. Black filled circles show stable equilibria, unfilled are unstable equilibria, partially filled is a saddle-node bifurcations. Crosses show saddle equilibria. Saddle-node-homoclinic cycles in (b), and limit cycles in (c) and (d) are in yellow.
III Excitable decision-making
We present the E-NOD model in Section III-A, and analyze its behavior through geometric phase-plane analysis in Section III-B.
III-A E-NOD for a Single Agent
We define E-NOD by introducing and coupling a slow regulation variable $u_{s}$ to NOD (2):
$\tau_{z}\dot{z}=-d\,z+\tanh\Big{(}(u_{0}-u_{s}+K_{u}z^{2})\cdot\left(\alpha z\right)+b\Big{)}$ (4a)
$\tau_{u_{s}}\dot{u}_{s}=K_{u_{s}}z^{4}-u_{s}$ (4b)
where $\tau_{u_{s}}\!\gg\!\tau_{z}$ is larger by at least an order of magnitude such that $u_{s}$ increases more slowly than opinion $z$. E-NOD (4) describes dynamics with excitability: a fast positive feedback (mediated by $z$) acts to excite the system, while a slow negative feedback (mediated by $u_{s}$) regulates it back to near its ultrasensitive pitchfork singularity.
The fast positive feedback in (4a) comes from the term $(u_{0}+K_{u}z^{2})(\alpha z)$, as in NOD (2). The slow negative feedback comes from the coupling of the new variable $u_{s}$. As $z$ grows in magnitude (whether positive or negative, i.e., independent of which option is preferred), $u_{s}$ also grows, although more slowly, according to (4b). When $u_{s}$ becomes large, because of the $-u_{s}(\alpha z)$ term in  (4a), it drives $z$ back towards zero. Then as $z$ decreases towards zero, $u_{s}$ also decreases back towards zero according to (4b). The result is a spike in both $z$ and $u_{s}$.
III-B Geometric Analysis of E-NOD
We use phase-plane analysis to study and illustrate the spiking and decision-making behavior of E-NOD (4). To construct the phase-plane, we first compute the nullclines.
The $z$-nullcline is defined as the solutions $z$ as a function of $u_{s}$ that satisfy $\dot{z}\!=\!0$ for (4a). This is the same as solving for the equilibrium solutions of (2) as a function of $u_{0}$, except now it is a function of $u_{s}$. Thus, the $z$-nullcline (pink line in Fig. 4) corresponds to the bifurcation diagram of (2), mirrored about the vertical axis and shifted right by $u_{0}$ Here, when $b\!=\!0$, the neutral solution $z\!=\!0$ is stable (unstable) for $u_{s}\!>\!u_{s}^{*}$ ($u_{s}\!<\!u_{s}^{*}$).
The $u_{s}$-nullcline (blue line in Fig. 4) is defined as the solutions that satisfy $\dot{u}_{s}\!=\!0$ in (4b), which gives the quartic function $u_{s}\!=\!K_{u_{s}}z^{4}$. This describes a quartic parabola with symmetric solutions about the $u_{s}$ axis and vertex at $u_{s}\!=\!0$. As $K_{u}$ increases, this parabola becomes narrower.
We next examine the intersections of the nullclines, which determine equilibrium solutions of E-NOD (4). As illustrated in Fig. 4 the intersections depend on the value of $u_{0}$. If $b\!=\!0$, the neutral state $(u_{s},z)\!=\!(0,0)$ is always an equlibrium and two more equilibria, symmetric about $z\!=\!0$, may be present for sufficiently high $u_{0}$.
Fig. 4a illustrates the phase-plane when $b\!=\!0$ and $u_{0}\!<\!u_{0}^{th}$, $u_{0}^{th}>0$ a threshold. The nullclines have only one point of intersection at the neutral state. The neutral state is stable. Trajectories will converge to and settle at this point and no excitable behavior in the decision-making will take place.
Fig. 4b illustrates the phase-plane when $b\!=\!0$ and $u_{0}\!=\!u_{0}^{th}$. The nullclines shown have three points of intersection, the neutral state and two unstable equilibria symmetric about $z\!=\!0$. The neutral state is a saddle-node bifurcation with one exponentially stable eigendirection (parallel to the $u_{s}$-axis $z=0$) and one marginally unstable direction (parallel to the $z$-axis $u_{s}=0$). There are two saddle-node-homoclinic (infinite period) cycles, one diverging upward from the saddle-node and another diverging downward from it. In the absence of noise/exogenous perturbations, all trajectories asymptotically converge to the saddle-node. The presence of any, arbitrarily small, noisy perturbation, makes the trajectories escape from the saddle-node at random time instants along either the upward or downward saddle-node-homoclinic cycle, leading to large prototypical excursions in the $(u_{s},z)$ plane
These large prototypical excursions are highly reminiscent of “spiking” trajectories of excitable neuronal system. By analogy, we call them “decision spikes” or “excitable decisions”. However, in contrast to neuronal spikes which happen only in one direction, decision spikes can happen in multiple directions, i.e., in as many directions as there are options. For the one-dimensional two options dynamics studied here, both upward (in favor of option 1) and downward (in favor of option 2) decision spikes are possible.
When $b\!=\!0$ and $u_{0}\!>\!u^{th}$ (Fig. 4c) the three fixed point are unstable and it is possible to prove, along the same lines as [20], the existence of two limit cycles, mirror symmetric about the horizontal axis $z=0$. These limit cycles are made of repetitive decision spikes, i.e., spiking decision limit cycles. Geometric singular perturbation analysis [21, 22] provides the tools to rigorously prove the existence of these spiking decision limit cycles. Such an analysis goes beyond the scope of this paper. Instead, we leverage Fig. 4c to describe qualitatively a typical oscillatory spiking decision behavior in the presence of small noisy perturbations.
Consider a trajectory with initial conditions in the rightmost part of the phase portrait. Initially, the trajectory rapidly converges to the $z=0$ axis and then slowly slides leftward approaching the neutral state. As soon as the trajectory hits this equilibrium, noisy perturbations push it either upward or downward, generating an upward or downward decision spike respectively. The decision spike trajectory brings the trajectory back to the pitchfork singularity, the next decision spike is generated, and the spiking decision cycle continues.
When input $b\!\neq\!0$ and $|b|$ is sufficiently large, the $z$-nullcline unfolds accordingly to the universal unfolding of the pitchfork [17, Ch. III]. Due to the nullcline unfolding, the phase-plane geometry changes qualitatively as shown in Fig. 4d. Similar geometric singular perturbation analysis methods as those employed for the analysis of Fig. 4b and 4c reveal the existence of a unique spiking decision limit cycle associated to spiking decisions toward the option favored by the inputs (option 1, upward decision spikes in the case of Fig. 4d where $b>0$ provides evidence in favor of option 1).
Observe that in the presence of informative inputs (Fig. 4d) the decision spiking frequency is higher than in the case of endogenous decision spiking oscillations (Fig. 4c). This feature is akin to spike frequency coding of input intensity in biological neural systems. In applications like robot navigation, $u_{0}$ can be controlled to avoid endogenous spiking.
IV Multi-agent excitable decision-making
It is possible to generalize the E-NOD model to the case of $N_{a}$ agents in a way analogous to how the NOD is generalized [1] but including the new slow regulation variable(s). There are two natural, non-equivalent, ways of doing this. One is to introduce a centralized regulation variable $u_{s}$, and the other is to introduce distributed regulation variables $u_{s,i}$. We focus on the distributed solution. For each $1\!\leq\!i\!\leq\!N_{a}$, agent $i$ has two state variables $z_{i}$ and $u_{s,i}$ with dynamics given by
$\textstyle\dot{z}_{i}=-d\,z_{i}+\tanh\left((u_{0}-u_{s,i}+K_{u}z_{i}^{2})\left(\sum_{k=1}^{N_{a}}a_{ik}z_{k}+b_{i}\right)\right)$ (5a)
$\dot{u}_{s,i}=-u_{s,i}+K_{u_{s}}z_{i}^{4}$ (5b)
where $A\!=\![a_{ik}]$ is the E-NOD network adjacency matrix, capturing both opinion self-reinforcing terms (diagonal elements) and inter-agent opinion exchange terms (out-diagonal elements). We assume homogeneous agents, i.e., all agents have the same $d$, $u_{0}$, $K_{u}$ and $K_{u_{s}}$. Notice that (5) is exactly the networked, distributed version of (4).
The higher dimensional nature of this system makes its dynamics harder to characterize analytically. However, many of its properties can be inferred from our analysis of the one dimensional case, the analysis of the NOD in [1], and/or from numerical simulations. Here we just list a few notable properties of (5) that appear to hold under mild assumptions. A rigorous analysis is left for future works.
The origin $\textbf{0}\!\in\!\mathbb{R}^{n}$ is always a fixed point for the no input ($b_{i}\!=\!0$) case, but its stability depends on the parameter choice. There is a bifurcation of the origin with respect to parameter $u_{0}$ where the origin goes from being stable to unstable at a critical value $u_{0}^{*}$. For appropriate parameters, one or two limit cycles, symmetric with respect to some hyperplane, appear at this bifurcation. These limit cycles have in general longer periods as compared to the single agent case with the same parameters. For values of $u_{0}$ slightly smaller than $u_{0}^{*}$ the system is excitable and all trajectories approach one of two prototypical spiking decision trajectories (i.e., upward or downward spikes), both of which converge back to the origin. At $u_{0}\!=\!u_{0}^{*}$ these two spiking decision trajectories bifurcate into one or two attracting limit cycles.
When the graph induced by $A$ is regular, then the consensus subspace, $\{(u_{s}\!\cdot\!\mathds{1},z\!\cdot\!\mathds{1})\!\mid\!u_{s},z\!\in\!\mathds{R}\}$ with $\mathds{1}\!=\!(1,\ldots,1)\!\in\!\mathds{R}^{N_{a}}$, is forward invariant. The dynamics on this subspace can be reduced to those of the single agent model. If this subspace is stable, the dynamics lead to opinion synchronization.
V Application to agile, social robot navigation
We use the E-NOD (4) to design an agile controller for a social robot navigating around oncoming human movers. Simulations of the resulting behavior can be seen in Fig. 1 and 5 and are animated here.
Consider a robot moving toward a goal at constant speed in a space that may be occupied by other robots and humans. Each object in the 2D space is described by its position and direction of motion. Each robot seeks to avoid collisions with humans and reach its goal. For each robot, we introduce an E-NOD state $z$ ruled by (4) that determines the robot preference for turning left ($z\!>\!0$), right ($z\!<\!0$), or going straight ($z\!=\!0$), relative to its nominal (goal-tracking) control, in order to avoid other movers and obstacles. That is, we embody the E-NOD (4) in the physical robot steering control. For multi-robot coordination, we use the multi-agent E-NOD (5).
For human-robot interaction and collision avoidance, we define a proxy of the human’s (unknown) heading opinion at time $t$, $\hat{z}_{h}(t)\!=\!\tan(\eta_{h}(t))$, where $\eta_{h}(t)$ is a reconstruction at time $t$ of the human heading direction from the robot viewpoint, i.e., the angle between the (measured) human’s heading direction and the direction of the robot-human vector. Each robot steering opinion $z_{i}$ evolves over time $t$ according to (5), where the opinion interaction weight $a_{ik}$ depends on the distance between robots $i$ and $k$. Human heading direction proxy is added as an input $b\!=\!u\!\cdot\!\hat{z}_{h}$. The basal attention of each robot $u_{0}$ is modulated by the proximity and bearing of a robot to a human. As a robot and human move closer to an impending collision, $u_{0}$ increases.
Fig. 1 compares the trajectory and opinion of one robot navigating around one human mover when controlled by NOD (adapted from [11]) or E-NOD (4). Initially the human appears to pass on the robot’s left, so both robots bear right (with $z\!<\!0$). When the human abruptly turns to pass on the robot’s right, NOD lags to adapt and bear left ($z\!>\!0$), whereas E-NOD allows for a quick and agile change in heading direction. Only the E-NOD controlled robot is able to turn left soon enough to navigate around the human mover.
Fig. 5 showcases a set of three robots coupled through multi-agent E-NOD (5) in a leader-follower configuration as they navigate toward a common target and around two approaching humans. The only difference between the simulation in Fig. 5a and Fig. 5b is the sign of the communication network edge weights. In Fig. 5a, observers follow leader opinion. When the leader robot turns to the left to dodge the approaching humans, all robots follow, gracefully reaching the goal. In Fig. 5b, observers do not follow leader opinion. When the leader robot turns left with $z\!>\!0$, the second robot disagrees and forms $z\!<\!0$ to turn right. In further disagreement, the last robot turns left with $z\!>\!0$. The second robot later interacts with another human in its path, and its opinion adapts to $z\!>\!0$, leading to a timely left turn to avoid a collision. These simulatons show the agility of E-NOD and its sequential decision-making power.
Figure 5: Trajectories of social robot leader-follower teams navigating around approaching human movers. (a) Observers follow leader opinion. (b) Observers do not follow leader opinion. Plot of $z$ and $u_{s}$ over time $t$ for each robot in this case.
VI Final Remarks
We presented Excitable Nonlinear Opinion Dynamics (E-NOD) for a single agent and two options and showcased its ability to swiftly form opinions and regulate back to ultrasensitivity. Parameter $K_{u}$ was identified as a cause for NOD becoming too robust, but the self-regulation of E-NOD recovers flexibility. The existence of limit cycles for certain parameter regimes in E-NOD was found and examined. We presented a multi-agent E-NOD model and highlighted its limit cycles and potential for agent synchronization. We illustrated the agility of E-NOD in an application to social robot navigation. We aim to provide analytical guarantees on the onset of periodic spiking in limit cycles. We will also implement E-NOD control on physical robots in social navigation experiments.